# “Fear will keep the local systems in line.”

### v2

@@ 8.9. Geometric Langlands for non-pure inner forms
$0\to\pi_{1}({\check{G}})\to\pi_{{\check{G}}_{\operatorname{ad}}}\to Z_{\check{G}}\to 0.$
@@
We expect that equivalences parallel to Corollaries 8.9.3 and 8.9.8 also take place in the framework of Fargues-Scholze theory of [FS].
8.10. Arithmetic consequences
@@
where we view ${\mathcal{G}}_{{\mathfrak{G}}_{Z_{G}}}$ as a $\mu_{\infty}(\overline{{\mathbb{F}}}_{q})$-gerbe on $\operatorname{Ge}_{\pi_{1}({\check{G}})}$, and we turn it into a ${\mathbb{G}}_{m}$-gerbe via
$\mu_{\infty}(\overline{{\mathbb{F}}}_{q})\hookrightarrow\mu_{\infty}(\overline{{\mathbb{Q}}}_{\ell})\subset{\mathbb{G}}_{m}.$
8.10.2.
@@
where we view ${\mathcal{G}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}}$ as a $\mu_{\infty}(\overline{{\mathbb{F}}}_{q})$-gerbe on $\operatorname{Ge}_{\pi_{1}(G)}$, and we turn it into a $\overline{{\mathbb{Q}}}_{\ell}^{\times}$-gerbe via
(8.44) $\mu_{\infty}(\overline{{\mathbb{F}}}_{q})\hookrightarrow\mu_{\infty}(\overline{{\mathbb{Q}}}_{\ell})\subset\overline{{\mathbb{Q}}}_{\ell}^{\times}.$
8.10.3.
@@
The $\overline{{\mathbb{Q}}}_{\ell}^{\times}$-gerbe ${\mathsf{p}}^{*}_{\pi_{1}(G)}({\mathcal{G}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}})$ (see (8.44)) on $\operatorname{Bun}_{G}$ gives rise to a $\overline{{\mathbb{Q}}}_{\ell}^{\times}$-torsor over $\operatorname{Bun}_{G}({\mathbb{F}}_{q})$, to be denoted ${\mathcal{P}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}}$. Consider the space
$\operatorname{Sect}_{c}(\operatorname{Bun}_{G}({\mathbb{F}}_{q}),{\mathcal{P}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}})$
of its compactly supported sections.
@@
(8.46) $\operatorname{Sect}_{c}(\operatorname{Bun}_{G}({\mathbb{F}}_{q}),{\mathcal{P}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}})\simeq\Gamma(\operatorname{LS}_{{\check{G}},{\mathfrak{G}}_{Z_{{\check{G}}}}}^{\operatorname{arithm}},\omega_{\operatorname{LS}_{\check{G}}^{\operatorname{arithm}}}).$

### v3

@@ 8.10. Geometric Langlands for non-pure inner forms
$0\to\pi_{1}({\check{G}})\to\pi_{1}({\check{G}}_{\operatorname{ad}})\to Z_{\check{G}}\to 0.$
@@
We expect that equivalences parallel to Corollaries 8.10.3 and 8.10.8 also take place in the framework of Fargues-Scholze theory of [FS].
8.11. An arithmetic variant
The contents of this subsection are not used elsewhere in the paper, and and be considered as a collection of informal remarks. Our goal here is to explain that Corollaries 8.10.3 and 8.10.8 are not just category-theoretic, but have meaningful implications for automorphic functions.
@@ 8.11. An arithmetic variant
where we view ${\mathcal{G}}_{{\mathfrak{G}}_{Z_{G}}}$ as naturally a $\mu_{\infty}(\overline{{\mathbb{Q}}}_{\ell})$-gerbe on $\operatorname{Ge}_{\pi_{1}({\check{G}})}$, and we turn it into a ${\mathbb{G}}_{m}$-gerbe via
$\mu_{\infty}(\overline{{\mathbb{Q}}}_{\ell})\subset{\mathbb{G}}_{m}.$
8.11.2.
@@
where we view ${\mathcal{G}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}}$ as a $\mu_{\infty}(\overline{{\mathbb{Q}}}_{\ell})$-gerbe on $\operatorname{Ge}_{\pi_{1}(G)}$.
8.11.3.
@@
The $\overline{{\mathbb{Q}}}_{\ell}^{\times}$-gerbe ${\mathsf{p}}^{*}_{\pi_{1}(G)}({\mathcal{G}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}})$ on $\operatorname{Bun}_{G}$ gives rise to a $\overline{{\mathbb{Q}}}_{\ell}^{\times}$-torsor over $\operatorname{Bun}_{G}({\mathbb{F}}_{q})$, to be denoted ${\mathcal{P}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}}$. Given this torsor, consider the corresponding space of twisted compactly-supported functions
$\operatorname{Funct}_{c}(\operatorname{Bun}_{G}({\mathbb{F}}_{q}))_{{\mathcal{P}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}}}:=\operatorname{Sect}_{c}(\operatorname{Bun}_{G}({\mathbb{F}}_{q}),{\mathcal{P}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}}\overset{\overline{{\mathbb{Q}}}_{\ell}^{\times}}{\times}\overline{{\mathbb{Q}}}_{\ell}),$
@@
(8.48) $\operatorname{Funct}_{c}(\operatorname{Bun}_{G}({\mathbb{F}}_{q}))_{{\mathcal{P}}_{{\mathfrak{G}}_{Z_{{\check{G}}}}}}\simeq\Gamma(\operatorname{LS}_{{\check{G}},{\mathfrak{G}}_{Z_{{\check{G}}}}}^{\operatorname{arithm}},\omega_{\operatorname{LS}_{\check{G}}^{\operatorname{arithm}}}).$